# Stochastic Subspace Descent Accelerated via Bi-fidelity Line Search

### v1

@@ Abstract
Efficient optimization remains a fundamental challenge across numerous scientific and engineering domains, especially when objective function and gradient evaluations are computationally expensive. While zeroth-order optimization methods offer effective approaches when gradients are inaccessible, their practical performance can be limited by the high cost associated with function queries. This work introduces the bi-fidelity stochastic subspace descent (BF-SSD) algorithm, a novel zeroth-order optimization method designed to reduce this computational burden. BF-SSD leverages a bi-fidelity framework, constructing a surrogate model from a combination of computationally inexpensive low-fidelity (LF) and accurate high-fidelity (HF) function evaluations. This surrogate model facilitates an efficient backtracking line search for step size selection, for which we provide theoretical convergence guarantees under standard assumptions. We perform a comprehensive empirical evaluation of BF-SSD across four distinct problems: a synthetic optimization benchmark, dual-form kernel ridge regression, black-box adversarial attacks on machine learning models, and transformer-based black-box language model fine-tuning. Numerical results demonstrate that BF-SSD consistently achieves superior optimization performance while requiring significantly fewer HF function evaluations compared to relevant baseline methods. This study highlights the efficacy of integrating bi-fidelity strategies within zeroth-order optimization, positioning BF-SSD as a promising and computationally efficient approach for tackling large-scale, high-dimensional problems encountered in various real-world applications.
@@ 1.1 Related Work
Multi-fidelity is a well-established concept in engineering and scientific computing for reducing computational costs. It has been widely applied across various domains, including aerodynamic design (Zhang et al., 2021), structural optimization (Ng & Willcox, 2014; De et al., 2020), data sampling (Cheng et al., 2024b), and uncertainty quantification (Peherstorfer et al., 2018; Cheng et al., 2024a; De & Doostan, 2022; De et al., 2023; Cheng & Doostan, 2025). As one sub-branch, multi-fidelity optimization has been employed in hyperparameter tuning (Wu et al., 2020), accelerating Bayesian optimization (Kandasamy et al., 2016; Takeno et al., 2020), and reinforcement learning (Cutler et al., 2014) within machine learning. However, despite its relevance in settings where function evaluations are costly, its application in zeroth-order optimization remains largely unexplored (de Montbrun & Gerchinovitz, 2024) and has not been applied to any randomized zeroth-order method.
@@ 1.2 Contributions
In this work, we propose a multi-fidelity line search scheme. Unlike previous works that use static surrogate (reduced-order) models which do not change from iteration to iteration (Yue & Meerbergen, 2013; Grundvig, 2023) , our method constructs a temporary surrogate after the gradient is estimated. This allows us to focus on building a one-dimensional surrogate which is a much easier task than building an accurate $D$-dimensional surrogate to $f$. By leveraging a LF model, we construct a simple linear surrogate using a small number of HF evaluations, $n_{k}$. This surrogate facilitates the identification of an optimal step size, assuming certain conditions between the LF and HF models hold.
@@
    • Black-box image-based adversarial attacks with a LF model trained via knowledge distillation;
@@ 2 Line Search on Bi-fidelity Surrogate
[…]
@@ 2.1 Algorithm
    Given the current position $\bm{x}_{k}\in{\mathbb{R}}^{D}$, gradient $\bm{v}_{k}\in{\mathbb{R}}^{D}$, and initial step size $\alpha_{\textup{max}}\in{\mathbb{R}}$, sample $n_{k}$ equi-spaced HF evaluations in $[0,\alpha_{\textup{max}}]$ and build the surrogate $\tilde{\varphi}_{k}:{\mathbb{R}}\to{\mathbb{R}}$ following Equation (1.3) (see Algo. 1 for details);
@@
[…]
@@ 2.2 Convergence Results
For each iteration $k$, the surrogate must be accurate; i.e., it satisfies Equation (2.8). In particular, a sufficient condition is that the number of HF evaluations, $n_{k}$, for building the surrogate $\tilde{\varphi}_{k}$ is such that
$n_{k}\geq\frac{WL(1+c)\alpha_{\textup{max}}}{c\beta\lVert\bm{v}_{k}\rVert^{2}},\quad\text{i.e.,}\quad n_{k}=\Omega\left(\frac{W(\alpha_{\textup{max}}L+1)}{\lVert\bm{v}_{k}\rVert^{2}}\right).$ (2.4)
@@
Given an initial point $\bm{x}_{0}$, the algorithm in Section 2.1 generates a sequence $(\bm{x}_{k})$ such that
@@
$N_{\epsilon}=\sum_{k=1}^{K_{\epsilon}}\left(n_{k}+\mathcal{O}(D)\right)=\mathcal{O}\left(\frac{WL^{2}}{\epsilon}+\frac{DL}{\epsilon}\right).$ (2.6)
The result in Equation (2.6) suggests that the number of function evaluations for $\epsilon$-convergence for the proposed algorithm can be significantly reduced when $W$ is small.
@@
The proof of Remark 2.6 follows the convergence proof of gradient descent using backtracking line search. Comparing the results in Equations (2.6) and (2.7), we observe that the advantage of using our bi-fidelity surrogate depends on the value of $W$. Notice that if $W$ is sufficiently small so that $WL^{2}\leq L\log L$, then the worst-case bound of our method is better than that of the zeroth-order gradient descent. We emphasize that our convergence result in Equation (2.6) is loose, due to the global nature of the Assumption 2.2 and difficulty in precisely describing the quality of the LF function relative to its HF counterpart. Hence, we view our convergence analysis as a reassurance that the method does converge, and rely on numerical experiments to elucidate when the method improves over baseline methods.
@@
When the step size satisfies $\alpha_{k}\in[c/(L+cL),1/(L+cL)]$, the quadratic inequality $-\alpha_{k}+\alpha_{k}^{2}L/2+c/(2(1+c)^{2}L)\leq-\alpha_{k}/2$ holds, along with the fact that $\beta\leq 1/2$, which implies the following bi-fidelity-adjusted Armijo condition
@@ 2.4 Examples of Possible Low-Fidelity Functions
The most ideal case occurs when the HF model is an affine transformation of the LF model, i.e., $f^{\textup{HF}}(\bm{x})=\rho f^{\textup{LF}}(\bm{x})+c$. In this case, the Lipschitz constant $W$, as defined in Assumption 2.2, is zero, and the number of function evaluations required for convergence is proportional to the number of iterations, as $n_{k}=1$ is sufficient.
@@ 3 Bi-Fidelity Line Search with Stochastic Subspace Descent
1:  Define $\{(\tilde{\alpha}_{j},\varphi(\tilde{\alpha}_{j}))\}_{j=0}^{n_{k}}$ as equispaced points between 00 and $\alpha_{\textup{max}}$ (including endpoints), and compute HF evaluations $\varphi(\tilde{\alpha}_{j})\leftarrow f^{\textup{HF}}(\bm{x}_{k}+\tilde{\alpha}_{j}\bm{v}_{k})$;
2:  $\rho_{k}\leftarrow f^{\textup{HF}}(\bm{x}_{k})/f^{\textup{LF}}(\bm{x}_{k})$;
3:  $\psi(\tilde{\alpha}_{j})\leftarrow\varphi(\tilde{\alpha}_{j})-\rho_{k}f^{\textup{LF}}(\bm{x}_{k}+\tilde{\alpha}_{j}\bm{v}_{k}),\quad j=1,\dots,n_{k}$;
@@
The proposed bi-fidelity line search algorithm, combined with SSD, will be referred to as bi-fidelity SSD (BF-SSD), and is summarized in Algorithm 3. Our theory covers either $\ell=D$ with bi-fidelity line search (Thm. 2.4) or $1\leq\ell\leq D$ with HF line search (Appendix B). Combining the two analyses is fairly complicated, and we defer it to a future study.
For practical purposes, since the parameters in Assumption 2.3 are often unknown, we set $\rho_{k}$ as described above and choose $n_{k}=1$ to minimize the cost of generating $\tilde{\varphi}_{k}(\alpha)$. This choice also demonstrates excellent empirical performance across all HF and LF pairs we have examined.
@@ 4 Empirical Experiments
The performance of the optimizers is assessed based on the number of HF objective function evaluations required, accounting for LF calls (in terms of fractional equivalent HF function calls) as appropriate.  [^1] The code is available at https://github.com/CU-UQ/BF-Stochastic-Subspace-Descent/.
@@ 4.1 Synthetic Problem: Worst Function in the World
where $x_{i}$ denotes the $i^{\textup{th}}$ entry of the input $\bm{x}$ and $r<D$ is a constant integer defining the intrinsic dimension of the problem. The function is convex with global minimum value 00. The Lipschitz constant of the gradient of this function is $L$. Nesterov (2013) has shown that a wide range of iterative first-order methods perform poorly when minimizing $f(\bm{x};r,L)$ with initial point $\bm{x}_{0}=\bm{0}$.
@@
We set the dimension $D=1,\!000$, $\ell=20$, and $L=20$. The intrinsic dimensions of the LF and HF functions are $r_{H}$ and $r_{L}$, respectively. We choose $r_{L}\ll r_{H}$ and assume the computational cost ratio between HF and LF evaluations is $r_{H}:r_{L}$. For Gradient Descent, we choose the standard step size of $0.051/L=$ 0.05, and for the GS and SSD-based methods, the step size is $\ell/(LD)$. The backtracking parameter is $\beta=\ell/(2D)$. The hyperparameter study is conducted according to different values of $c\in\{0.8,0.9,0.99\}$ and $\ell\in\{5,10,20\}$. All the experiments are repeated 10 times, with shaded regions denoting the worst and the best performance over 10 trials.
@@
Equivalent HF function evaluations $N$
@@
FS-SSD $2.47\pm 0.00$ $2.45\pm 0.00$ $2.26\pm 0.00$ $2.08\pm 0.00$ $1.92\pm 0.00$
SPSA $1.95\pm 0.05$ $\mathbf{0.61\pm 0.04}$ $0.19\pm 0.00$ $0.14\pm 0.00$ $0.12\pm 0.00$
GS $2.47\pm 0.00$ $2.46\pm 0.00$ $2.36\pm 0.00$ $2.25\pm 0.00$ $2.15\pm 0.00$
HF-SSD $2.07\pm 0.15$ $0.77\pm 0.06$ $0.22\pm 0.02$ $0.17\pm 0.02$ $0.14\pm 0.02$
BF-SSD $2.00\pm 0.05$ $0.66\pm 0.03$ $\mathbf{0.17\pm 0.00}$ $\mathbf{0.11\pm 0.00}$ $\mathbf{0.09\pm 0.00}$
VR-SSD $2.47\pm 0.00$ $2.10\pm 0.02$ $0.63\pm 0.01$ $0.43\pm 0.00$ $0.35\pm 0.00$
Table 1: Performance values (mean $\pm$ std over 10 runs) showing the objective function for different optimization methods at various HF function evaluations $N$ with $\ell=20$ and $c=0.99$. The minimum values in each column are highlighted in bold.
@@
FS-SSD $2.0766\pm 0.0038$ $1.6726\pm 0.0057$ $1.2943\pm 0.0047$ $\mathbf{0.9447\pm 0.0041}$
HF-SSD $0.1745\pm 0.0209$ $\mathbf{0.1357\pm 0.0073}$ $0.1482\pm 0.0067$ $0.1893\pm 0.0115$
BF-SSD $\mathbf{0.1149\pm 0.0016}$ $0.1206\pm 0.0015$ $0.1236\pm 0.0028$ $0.1329\pm 0.0024$
VR-SSD $0.4328\pm 0.0030$ $0.4268\pm 0.0023$ $0.4232\pm 0.0018$ $\mathbf{0.4227\pm 0.0016}$

### v2

@@ Abstract
Efficient optimization remains a fundamental challenge across numerous scientific and engineering domains, particularly when objective function evaluations are computationally expensive and gradient information is inaccessible. While zeroth-order optimization methods address the lack of gradients, their performance often suffers due to the high cost of repeated function queries. This work introduces a bi-fidelity line search scheme tailored for zeroth-order optimization. Our method constructs a temporary surrogate model by strategically combining inexpensive low-fidelity (LF) evaluations with accurate high-fidelity (HF) evaluations of the objective function. This surrogate enables an efficient backtracking line search for step size selection, significantly reducing the number of costly HF queries required. We provide theoretical convergence guarantees for this scheme under standard assumptions. Furthermore, we integrate this bi-fidelity strategy into the stochastic subspace descent framework, proposing the bi-fidelity stochastic subspace descent (BF-SSD) algorithm. A comprehensive empirical evaluation of BF-SSD is conducted across four distinct problems: a synthetic optimization benchmark, dual-form kernel ridge regression, black-box adversarial attacks on machine learning models, and transformer-based black-box language model fine-tuning. The numerical results consistently demonstrate that BF-SSD achieves superior optimization performance, particularly in terms of solution quality obtained per HF function evaluation, when compared against relevant baseline methods. This study highlights the efficacy of integrating bi-fidelity strategies within zeroth-order optimization frameworks, positioning BF-SSD as a promising and computationally efficient approach for tackling large-scale, high-dimensional optimization problems encountered in various real-world applications.
@@ 1.1 Related Work
Multi-fidelity is a well-established concept in engineering and scientific computing for reducing computational costs. It has been widely applied across various domains, including aerodynamic design (Zhang et al., 2021), structural optimization (Ng & Willcox, 2014; De et al., 2020), data sampling (Cheng et al., 2024b), and uncertainty quantification (Peherstorfer et al., 2018; Cheng et al., 2024a; De & Doostan, 2022; De et al., 2023; Cheng & Doostan, 2025; Cheng, 2025). As one sub-branch, multi-fidelity optimization has been employed in hyperparameter tuning (Wu et al., 2020), accelerating Bayesian optimization (Kandasamy et al., 2016; Takeno et al., 2020), and reinforcement learning (Cutler et al., 2014) within machine learning. However, despite its relevance in settings where function evaluations are costly, its application in zeroth-order optimization remains largely unexplored (de Montbrun & Gerchinovitz, 2024) and has not been applied to any randomized zeroth-order method.
@@ 1.2 Contributions
In this work, we propose a multi-fidelity line-search scheme. Unlike previous approaches that utilize static surrogate (e.g., reduced-order) models, which remain fixed throughout the optimization process (Yue & Meerbergen, 2013; Grundvig, 2023), our method constructs a temporary surrogate model in each iteration, specifically after the gradient (or search direction) has been estimated. This allows us to focus on building a one-dimensional surrogate along the search direction, a significantly simpler task compared to constructing expensive $D$-dimensional surrogates for the objective function $f$. By leveraging a computationally cheaper LF model, we construct a simple, local, linear surrogate using only a small number, $n_{k}$, of HF evaluations per iteration. Assuming certain conditions between the LF and HF models hold, this 1D surrogate is then used to efficiently identify a suitable step size.
@@
    • Black-box image-based adversarial attacks with an LF model trained via knowledge distillation;
@@ 2 Line Search on Bi-fidelity Surrogate
In this section, we discuss the proposed algorithm and present the main theoretical results derived in this work. Unless specified otherwise, $\lVert\cdot\rVert$ denotes the Euclidean norm for vectors and the spectral norm (i.e., the induced 2-norm) for matrices. For simplicity and to maintain focus on our primary contributions during the theoretical analysis, we assume in the proofs that $\bm{v}_{k}$ provides an accurate estimate of the high-fidelity gradient, specifically $\bm{v}_{k}\approx\nabla f^{\textup{HF}}(\bm{x}_{k})$. This assumption is employed solely for the theoretical development presented herein and does not hold for the practical implementation discussed in Section 3 or the numerical results presented in Section 4.
@@ 2.1 Algorithm
    Given the current position $\bm{x}_{k}\in{\mathbb{R}}^{D}$, gradient $\bm{v}_{k}\in{\mathbb{R}}^{D}$, and initial step size $\alpha_{\textup{max}}\in{\mathbb{R}}$, sample $n_{k}$ equi-spaced HF evaluations in $0,\alpha_{\textup{max}}]$ and build the surrogate $\tilde{\varphi}_{k}:{\mathbb{R}}\to{\mathbb{R}}$ following Equation (1.3) (see Algo. 1 for details);
@@
    $\text{s.t.}\quad\tilde{\varphi}_{k}(c^{m}\alpha_{\textup{max}};n_{k})$ $\leq f^{\textup{HF}}(\bm{x}_{k})-\beta c^{m}\alpha_{\textup{max}}\lVert\bm{v}_{k}\rVert^{2}.$
@@ 2.2 Convergence Results
Our next assumption, Assumption 2.3, is a sufficient condition that will be used in Lemma 2.7 to show that the surrogate is accurate.
@@
For each iteration $k$, we assume the number of HF evaluations, $n_{k}$, for building the surrogate $\tilde{\varphi}_{k}$ is sufficiently large such that
$n_{k}\geq\frac{WL(1+c)\alpha_{\textup{max}}}{c\beta\lVert\bm{v}_{k}\rVert^{2}},\quad\text{i.e.,}\quad n_{k}=\Omega\left(\frac{WL}{\lVert\bm{v}_{k}\rVert^{2}}\right),$ (2.4)
where $\Omega(\cdot)$ denotes a lower bound up to a constant difference.
@@
Given an initial point $\bm{x}_{0}$, assuming actual gradients are accurately estimated, the algorithm in Section 2.1 generates a sequence $(\bm{x}_{k})$ such that
@@
$N_{\epsilon}=\sum_{k=1}^{K_{\epsilon}}\left(n_{k}+\mathcal{O}(D)\right)=\mathcal{O}\left(\frac{WL^{2}}{\epsilon^{2}}+\frac{DL}{\epsilon}\right).$ (2.6)
@@
The proof of Remark 2.6 follows the convergence proof of gradient descent using backtracking line search. Comparing the results in Equations (2.6) and (2.7), we observe that the advantage of using our bi-fidelity surrogate depends on the value of $W$. Notice that if $W$ is sufficiently small so that $WL\leq\epsilon\log L$, then the worst-case bound of our method is better than that of the zeroth-order gradient descent. We emphasize that our convergence result in Equation (2.6) is loose, due to the global nature of the Assumption 2.2 and difficulty in precisely describing the quality of the LF function relative to its HF counterpart. Hence, we view our convergence analysis as a reassurance that the method does converge, and rely on numerical experiments to elucidate when the method improves over baseline methods.
@@
When the step size satisfies $\alpha_{k}\in[c/(L+cL),1/(L+cL)]$, the quadratic inequality $\alpha_{k}+\alpha_{k}^{2}L/2+c/(2(1+c)^{2}L)\leq-\alpha_{k}/2$ holds, along with the fact that $\beta\leq 1/2$, which implies the following bi-fidelity-adjusted Armijo condition
@@ 2.4 Examples of Possible Low-Fidelity Functions
The ideal case occurs when the HF model is an affine transformation of the LF model, i.e., $f^{\textup{HF}}(\bm{x})=\rho f^{\textup{LF}}(\bm{x})+c$. In this case, the Lipschitz constant $W$, as defined in Assumption 2.2, is zero, and the number of function evaluations required for convergence is proportional to the number of iterations, as $n_{k}=1$ is sufficient.
@@
This use-case satisfies the assumptions mentioned in Section 2.2. Assumption 2.1 is satisfied since the minimum is achieved ($f^{\textup{HF}}$ is continuous and coercive) and the Lipschitz constant of the gradient is $L=\|\bm{A}\|$. Assumption 2.2 is satisfied using the value of $W$ above, and Assumption 2.3 can be satisfied since it is just a parameter choice.
@@
This use-case also satisfies the assumptions mentioned in Section 2.2 under reasonable conditions. If each $\nabla f_{i}$ is Lipschitz continuous with constant $L_{i}$ (i.e., this is always true if $f_{i}$ is continuous and $\bm{x}$ is constrained to a compact set), then $\nabla f^{\textup{HF}}$ is $L$ Lipschitz continuous with $L=\frac{1}{n}\sum_{i=1}^{n}L_{i}$ via the triangle inequality, so under the mild assumption that the minimum is achieved, Assumption 2.1 is satisfied. Furthermore, Assumption 2.2 is satisfied using the value of $W$ above, and Assumption 2.3 can again be automatically satisfied since it is just a parameter choice.
@@ 3 Bi-Fidelity Line Search with Stochastic Subspace Descent
1: Define $\{(\tilde{\alpha}_{j},\varphi(\tilde{\alpha}_{j}))\}_{j=0}^{n_{k}}$ as equispaced points between 0 and $\alpha_{\textup{max}}$ (including endpoints), and compute HF evaluations $\varphi(\tilde{\alpha}_{j})\leftarrow f^{\textup{HF}}(\bm{x}_{k}+\tilde{\alpha}_{j}\bm{v}_{k})$;
[…]
@@
The proposed bi-fidelity line search algorithm, combined with SSD, will be referred to as bi-fidelity SSD (BF-SSD), and is summarized in Algorithm 3. Our theory covers either $\ell=D$ with bi-fidelity line search (Thm. 2.4) or $1\leq\ell\leq D$ with HF line search (Appendix B) under strongly convex, convex or non-convex settings. Combining the two analyses is fairly complicated, and we defer it to a future study. For reference, the bi-fidelity result is in Theorem 2.4 and below is one of the results from Appendix B.3 for the $1\leq\ell\leq D$ with HF line search case, in particular, the result covering the non-convex case:
Theorem 3.1 (proof in Appendix B.3).
$\min_{k\in\{0,\dots,K\}}\mathbb{E}[\lVert\nabla f(\bm{x}_{k})\rVert^{2}]\leq\max\left\{\frac{(f(\bm{x}_{0})-f^{*})}{(K+1)\beta\alpha_{\textup{max}}},\frac{DL(f(\bm{x}_{0})-f^{*})}{(K+1)\ell c\beta}\right\}.$ (3.4)
For practical purposes, since the parameters in Assumption 2.3 are often unknown, we set $\rho_{k}$ as described above and choose $n_{k}=1$ to minimize the cost of generating $\tilde{\varphi}_{k}(\alpha)$. This choice also demonstrates excellent empirical performance across all HF and LF pairs we have examined. We leave it as future work for adaptively choosing $n_{k}$, but suggest one possible scheme inspired by similar methods used in trust-region algorithms: starting with a small $n_{k}$, we find the stepsize via line search on the surrogate, as usual, and then compare the actual function value of the high-fidelity function with the predicted value from the surrogate. If the predicted decrease from the surrogate was overly optimistic, we then rebuild the surrogate with a larger $n_{k}$ (re-using existing samples) and repeat.
@@ 4 Empirical Experiments
  • Nesterov accelerated gradient descent (NAG): An accelerated zeroth-order gradient descent method by Nesterov that incorporates a decaying momentum term, which often leads to faster convergence. The gradient is estimated using forward differences.
@@
The performance of the optimizers is assessed based on the number of HF objective function evaluations required, accounting for LF calls (in terms of fractional equivalent HF function calls) as appropriate. While we also measured wall-clock time, the results strongly aligned with the equivalent HF call data. Therefore, for clarity, we present performance primarily in terms of equivalent HF calls.  [^1] The code is available at https://github.com/CU-UQ/BF-Stochastic-Subspace-Descent/.
@@ 4.1 Synthetic Problem: Worst Function in the World
where $x_{i}$ denotes the $i^{\textup{th}}$ entry of the input $\bm{x}$ and $r<D$ is a constant integer defining the intrinsic dimension of the problem. The function is convex with the global minimum value 0. The Lipschitz constant of the gradient of this function is $L$. Nesterov (2013) has shown that a wide range of iterative first-order methods perform poorly when minimizing $f(\bm{x};r,L)$ with initial point $\bm{x}_{0}=\bm{0}$.
@@
We set the dimension D=1, [^0] D=1,\!000, $\ell=20$, and $L=20$. The intrinsic dimensions of the LF and HF functions are $r_{H}$ and $r_{L}$, respectively. We choose $r_{L}\ll r_{H}$ and assume the computational cost ratio between HF and LF evaluations is $r_{H}:r_{L}$. For Gradient Descent, we choose the standard step size of $1/L=0.05$, and for the GS and SSD-based methods, the step size is $\ell/(LD)$. The backtracking parameter is $\beta=\ell/(2D)$. The hyperparameter study is conducted according to different values of $c\in\{0.8,0.9,0.99\}$ and $\ell\in\{5,10,20\}$. All the experiments are repeated 10 times, with shaded regions denoting the worst and the best performance over 10 trials.
@@
Equivalent HF function evaluations NN
@@
NAG $2.48\pm 0.00$ $2.48\pm 0.00$ $0.33\pm 0.00$ $6\pm 0.00 0.11±0$$1\pm 0.00$
FS-SSD $2.47\pm 0.00$ $45\pm 0.00 2.2$$26\pm 0.00$ $2.08\pm 0.00$ $1.93\pm 0.00$
SPSA $1.93\pm 0.00$ $\mathbf{0.66\pm 0.00}$ $0.20\pm 0.00$ $0.15\pm 0.00$ $0.12\pm 0.00$
GS $2.47\pm 0.00$ $2.46\pm 0.00$ $36\pm 0.00 2.2$$25\pm 0.00$ $2.15\pm 0.00$
HF-SSD $1.99\pm 0.05$ $0.75\pm 0.01$ $0.40\pm 0.21$ $0.37\pm 0.21$ $0.20\pm 0.06$
BF-SSD $2.00\pm 0.08$ $0.68\pm 0.02$ $\mathbf{0.17\pm 0.00}$ $\mathbf{0.11\pm 0.00}$ $\mathbf{0.09\pm 0.00}$
VR-SSD $2.47\pm 0.00$ $2.09\pm 0.01$ $0.62\pm 0.00$ $0.43\pm 0.00$ $0.35\pm 0.00$
Table 1: Performance values (mean $\pm$ std over 10 runs) showing the objective function for different optimization methods at various HF function evaluations $N$ with $\ell=20$ and $c=0.9$. The minimum values in each column are highlighted in bold.
@@
FS-SSD $2.0787\pm 0.0013$ $1.6621\pm 0.0003$ $1.2942\pm 0.0022$ $\mathbf{0.9473\pm 0.0027}$
HF-SSD $0.3696\pm 0.2109$ $\mathbf{0.1482\pm 0.0098}$ $0.1574\pm 0.0157$ $0.3696\pm 0.2368$
BF-SSD $\mathbf{0.1143\pm 0.0005}$ $0.1226\pm 0.0015$ $0.1260\pm 0.0032$ $0.1298\pm 0.0013$
VR-SSD $0.4309\pm 0.0025$ $0.4281\pm 0.0013$ $0.4238\pm 0.0009$ $\mathbf{0.4226\pm 0.0013}$